An ice chest contains 5 cans of apple juice, 6 cans of grape juice, 5 cans of orange juice and 4 cans of mango juice. Suppose that you reach into the container and randomly select three cans in succession. Find the probability of selecting 3 cans of apple juice.

Answer by Alan3354(69443)

An ice chest contains 5 cans of apple juice, 6 cans of grape juice, 5 cans of orange juice and 4 cans of mango juice. Suppose that you reach into the container and randomly select three cans in succession. Find the probability of selecting 3 cans of apple juice.
------------------------
Total of 5+6+5+4 cans = 20 cans
---
Without replacement:
1st choice is 5/20 = 1/4
2nd choice is 4/19
3rd choice is 3/18 = 1/6
Product = 1/114